# Probing flavored regimes of leptogenesis with gravitational waves from cosmic strings

### v1

@@ Abstract
Cosmic strings radiate detectable gravitational waves in models featuring high-scale symmetry breaking, e.g., high-scale leptogenesis. In this Letter, for the first time, we show that different flavored regimes of high-scale leptogenesis can be tested distinctly with the spectral features in cosmic string-radiated gravitational waves. This is possible if the scalar field that makes right-handed neutrinos massive is feebly coupled to the Standard Model Higgs. Each flavored regime, sensitive to low-energy neutrino experiments, leaves a marked imprint on the gravitational waves spectrum. A three-flavor and a two-flavor regime could be probed by a characteristic fall-off of the gravitational wave spectrum at the LISA-DECIGO-ET frequency bands with preceding scale-invariant amplitudes bounded from above and below. We present Gravitational Waves windows for Flavored Regimes of Leptogenesis (GWFRL) falsifiable in the upcoming experiments.
@@
In this Letter, we show that the $M_{N}$-dependent GW spectrum and a concrete link between GWs and high-scale leptogenesis can be achieved in the context of $U(1)_{B-L}$ model with a peculiar scalar field dynamics [25]. After the phase transition, $\Phi$ rolls down to its true vacuum and oscillates coherently to provide an early matter-domination when it is long-lived. If the lifetime of the scalar field gets determined by $v_{\Phi}$ and $f_{N}$, then $M_{N}$ regulates the beginning and the end of the matter-domination. Differently from Ref. [12], the amplitude and the spectral features of the GWs radiated from cosmic strings are imprinted by the $M_{N}$-dependent matter domination, thus carrying information on $M_{N}$, and therefore, different regimes of flavored leptogenesis. In particular, the scale invariance in the GW spectrum at higher frequencies breaks at a frequency $f_{\rm dec}$ determined by the end of the matter domination at the temperature $T_{\rm dec}(M_{N})$ (see Fig. 1). We point out that such a scenario allows us to probe the RHN mass scale markedly, in agreement with all the relevant constraints on the parameter space (see Fig. 2): i) the bound from the recent observations of nHz GWs by Pulsar Timing Arrays (PTA) [26, 27, 28, 29, 30] which disfavor GWs from cosmic strings [31]; ii) the condition that the cosmic strings radiate predominantly GWs than particles [32, 33]; iii) the requirement that the $\Phi$ decays before the Big-Bang-Nucleosynthesis (BBN) [34, 35]. We highlight that if the break of the GW spectrum occurs in the LISA (ET) frequency band, most likely it is a signature of a 3FL (2FL) regime. Remarkably, the amplitudes and the spectral break frequencies of the GW spectra predicted by our model are very well delimited and they are all within reach of planned gravitational wave experiments such as LISA, DECIGO and ET [36, 37, 38]. This ensures unique and highly testable GW signatures of flavor regimes of high-scale leptogenesis.
@@
The parameter space for $M_{N}$-dependent matter domination. The seesaw Lagrangian for neutrino masses and leptogenesis is given by
@@
where $f_{D}$ is the neutrino Dirac Yukawa coupling and $H(L)$ is the SM Higgs (lepton doublet). The fields $N_{R}$ and $L$ have $B-L$ charge $-1$, whereas the same charges for $\Phi$ and $H$ are $2$ and 00, respectively. The scalar field dynamics is dictated by the finite temperature potential restoring symmetry at higher temperatures given by [39, 40]
@@
where $D,E$ and $T_{0}$ are functions of $B-L$ gauge coupling $g^{\prime}$, $\lambda$ is the self-interaction coupling, and the zero temperature potential $V(\Phi,0)=-\frac{\mu^{2}}{2}\Phi^{2}+\frac{\lambda}{4}\Phi^{4}$ determines the vacuum expectation value $v_{\Phi}=\mu/\sqrt{\lambda}$. The last term in Eq. (2) generates a potential barrier causing a secondary minimum at $\Phi\neq 0$, which becomes degenerate with the $\Phi=0$ one at $T=T_{c}$. The potential barrier vanishes at $T_{0}~{}(\lesssim T_{c})$, and the minimum at $\Phi=0$ becomes a maximum. The transition from $\Phi=0$ to $\Phi=v_{\Phi}$ can be treated as a second-order transition with an extremely quick disappearance of the potential barrier if the order parameter $\Phi_{c}/T_{c}\ll 1$ (weakly first-order transition). Such a smooth transition can be obtained approximately within the ballpark of $\lambda\simeq{g^{\prime}}^{3}$ and for ${g^{\prime}}\lesssim 10^{-2}$ [13, 41]. Once $\Phi$ rolls down, it oscillates around $v_{\Phi}$ behaving as matter [42, 43, 25]. Three important observations are in order. First, the $U(1)_{B-L}$ breaking scale can be as high as the GUT scale, implying $v_{\phi}\lesssim 5\times 10^{15}~{}{\rm GeV}$. Second, the Hubble friction at $T_{c}$ should also be negligible for the rolling of the field down to $v_{\Phi}$; this can be achieved by considering $m_{\Phi}$ larger than the Hubble parameter $\mathcal{H}(T_{c})$. Third, RHNs become massive after the phase transition at $T=T_{c}$. Therefore, the scale of leptogenesis $T_{\rm lepto}\sim M_{N}$ is bounded from above as $T_{\rm lepto}\sim M_{N}\lesssim T_{c}$.
@@
We aim to make $\Phi$ long-lived to induce a matter-domination epoch with a lifetime determined by $v_{\phi}$ and $f_{N}$. In high-scale leptogenesis scenario, the coupling $f_{N}$ should be generically large to have $M_{N}$ comparable with $v_{\Phi}$. In such a case, if $m_{\Phi}>2M_{N}$, tree-level decay of $\Phi$ to RHN pairs ($\Phi\rightarrow NN$) would be too quick to make $\Phi$ long-lived. This restricts ourselves to the case $M_{N}>m_{\Phi}=\sqrt{2\lambda}v_{\Phi}$. Therefore, along with $M_{N}\lesssim T_{c}$, the $\Phi$ scalar is long-lived within the folowing range
$\sqrt{2g^{\prime}}g^{\prime}v_{\Phi}<M_{N}\lesssim 2\sqrt{g^{\prime}}v_{\Phi}\,,$ (3)
@@
where we have assumed $\lambda={g^{\prime}}^{3}$. In this case, the mass of $\Phi$ is suppressed by a factor $\sqrt{g^{\prime}}$ compared to the $M_{Z^{\prime}}=\sqrt{2}g^{\prime}v_{\Phi}$, thus kinematically forbidding the tree-level decays of $\Phi$ to $Z^{\prime}$ pairs. Note interestingly that the seesaw Lagrangian in Eq. (1) can provide an effective coupling $\lambda_{\Phi H,{\rm 1-loop}}\sim f_{D}^{2}f_{N}^{2}/(2\pi^{2})$ at one-loop to trigger $\Phi\rightarrow HH$ decay. Therefore, despite $f_{N}$ being large, we can have a smaller decay rate as it is proportional to $\lambda_{\Phi H,{\rm 1-loop}}^{2}$. This decay involves both: the RHN mass and the active neutrino mass $m_{\nu}$ via the Dirac Yukawa coupling $f_{D}=\sqrt{m_{\nu}f_{N}v_{\Phi}/v_{H}^{2}}$. In our discussion, we shall consider this radiative $\Phi\rightarrow HH$ decay with the assumption that $\lambda_{\Phi H,{\rm 1-loop}}\gg\lambda_{\Phi H,{\rm tree}}$. The decay rate up to a logarithmic factor is given by [44, 45, 25]
$\Gamma_{\Phi}^{HH}\simeq\Gamma_{0}\left(\frac{g^{\prime}}{10^{-5}}\right)^{3}\left(\frac{v_{\Phi}}{\rm 10^{13}~{}GeV}\right)^{2}\left(\frac{m_{\Phi}}{\rm 10^{8}~{}GeV}\right)\,,$ (4)
with $\Gamma_{0}=1.3\times 10^{-17}~{}{\rm GeV}$. We have assumed $m_{\nu}\simeq 0.01$ eV, the SM Higgs vacuum expectation value $v_{H}=174$ GeV, and $f_{N}\simeq{g^{\prime}}\ll 1$, which implies $M_{N}\simeq M_{Z^{\prime}}\ll v_{\Phi}$. The latter assumption should not drastically change the phenomenological implication of the scenario because of the constraint in Eq. (3).
The evolution of the energy density of $\Phi$ and the duration of matter domination epoch, which ends when $\Phi$ decay [25], can be tracked by solving the Friedman equations (see Supplemental Material I for further details). The decay temperature of the scalar field can be obtained analytically considering $\Gamma_{\Phi}^{HH}\simeq\mathcal{H}$, and is given by
$T_{\rm dec}\simeq 6.35\left(\frac{g^{\prime}}{10^{-5}}\right)^{3/4}\left(\frac{M_{N}}{10^{9}~{}{\rm GeV}}\right)^{3/2}{\rm GeV}\,,$ (5)
@@
where we consider $M_{N}=g^{\prime}v_{\Phi}$. The scalar decays must occur before BBN, i.e., $T_{\rm dec}\gtrsim T_{\rm BBN}\sim 10~{}{\rm MeV}$. Moreover, because the scalar field is long-lived, it produces entropy which dilutes any pre-existing relic, including the baryon asymmetry. The amount of entropy production in our scenario can be computed as $\kappa\simeq 140({g^{\prime}}/{10^{-3}})^{-1/4}({M_{N}}/{10^{9}~{}\rm GeV})^{-1/2}$. The entropy dilution by a factor of $\kappa^{-1}$ must be considered in the computation of baryon asymmetry. Even in the low $M_{N}$ region, the scenario does not allow large entropy dilution [25], consequently the baryon asymmetry can be comfortably generated. Finally, we mention that we can safely neglect the additional one-loop decay $\Phi\rightarrow f\bar{f}V$ (the two-body decay $\Phi\rightarrow f\bar{f}$ is suppressed due to chirality flip [46]), where $f$ and $V$ are SM fermions and vector bosons. Indeed, this decay dominates over the $\Phi\rightarrow HH$ only for $g^{\prime}\gtrsim 10^{-2}$ and smaller values of $v_{\Phi}$, which is outside our parameter space.
@@
Imprints of flavored leptogenesis on GWs from cosmic string. Gravitational waves are radiated from cosmic string loops chopped off from the long strings resulting from the spontaneous breaking of the gauged $U(1)_{B-L}$ [23, 24]. Long strings are described by a correlation length $L=\sqrt{\mu/\rho_{\infty}}$, with $\rho_{\infty}$ is the long string energy density and $\mu$ is the string tension defined as [47]
@@
$\Omega_{GW}(f)=\sum_{k=1}^{k_{\rm max}}\frac{2k\mathcal{F}_{\alpha}G\mu^{2}\Gamma_{k}}{f\rho_{c}}\int_{t_{i}}^{t_{0}}\left[\frac{a(t)}{a(t_{0})}\right]^{5}n_{\omega}\left(t,l_{k}\right){\rm d}t\,,$ (7)
@@
where $\beta=2/3(1+\omega)$ with $\omega$ being the equation of state parameter of the universe, and $A_{\beta}=5.4$ ($A_{\beta}=0.39$) for radiation-dominated (matter-dominated) universe [53]. The quantity $\Gamma_{k}={\Gamma k^{-\delta}}/{\zeta(\delta)}$ quantifies the emitted power in $k$-th mode, with $\delta=4/3$ ($\delta=5/3$) for loops containing cusps (kinks) [54]. We shall present the results for the dominant $k=1$ mode which captures the qualitative features of our analysis.
The integral in Eq. (7) is subjected to two Heaviside functions $\Theta(t_{i}-t_{\rm fric})\Theta(t_{i}-l_{c}/\alpha)$ which set cut-offs on the GW spectrum at very high frequencies $f_{*}$ above which it falls as $f^{-1}$. The quantity $t_{\rm fric}$ represents the time until which the motion of the string network is damped by friction [55], and $l_{c}$ is a critical length above which GW emission is dominant over particle production as shown by high-resolution numerical simulations [32, 33]. The critical length is $l_{c}\simeq\delta_{w}(\Gamma G\mu)^{-\gamma}$ where $\delta_{w}=(\sqrt{\lambda}v_{\Phi})^{-1}$ is the width of the string and $\gamma=2$ ($\gamma=1$) for loops containing cusps (kinks). Note that we deal with small values of $\lambda$ leading to “fat” strings, which are not conventional while studying GWs signatures from cosmic strings in BSM models [12]. In particular, in this case, the particle production cut-off in the GW spectrum is much stronger than the one usually considered for “thin strings” ($\lambda\simeq 1$). Thus, our scenario is also the first concrete BSM model that discusses the results of [32, 33] for small $\lambda$.
@@
Figure 1: The GW spectrum from cosmic string as predicted by our model for three benchmark choices for the parameter $g^{\prime}$ and $T_{\rm dec}$, with $M_{N}$ uniquely fixed by Eq. (5) (see text for details). Blue and green spectra correspond to 3FL, while the red one is for 2FL. The symbols show the analytical calculations of the GW plateau amplitude and the break frequency according to Eq.s (9) and (10), respectively.
@@
In Fig. 1, we show the GW spectrum from Eq. (7) (for $k=1$) predicted by our model for three benchmark scenarios: $g^{\prime}=10^{-6}$ and $T_{\rm dec}=0.05~{}{\rm GeV}$ (blue line), $g^{\prime}=10^{-5}$ and $T_{\rm dec}=1~{}{\rm GeV}$ (green line), $g^{\prime}=10^{-4}$ and $T_{\rm dec}=10^{3}~{}{\rm GeV}$ (red line), with the scale $M_{N}$ of leptogenesis being uniquely defined by Eq. (5) to be $1.3\times 10^{8}$, $2.9\times 10^{8}$ and $9.2\times 10^{9}$ in units of GeV, respectively. Without any intermediate matter-dominated epoch, the GW spectrum has two main contributions: a low-frequency peak owing to the GW radiation from the loops which are created in the radiation epoch and decay in the standard matter epoch, and an almost scale-invariant plateau at high frequencies as [48, 49, 56]
@@
$\Omega_{GW}^{\rm plt}(f)=\frac{128\pi\mathcal{F}_{\alpha}G\mu}{9\zeta(\delta)}\frac{A_{R}}{\epsilon_{R}}\Omega_{R}\left[(1+\epsilon_{R})^{3/2}-1\right]\,,$ (9)
@@
that arises from loop creation and decay in the radiation-dominated epoch only. In Eq. (9), $\epsilon_{R}=\alpha/\Gamma G\mu\gg 1$, $A_{R}\equiv A_{\beta}\simeq 5.4$ and $\Omega_{R}\sim 9\times 10^{-5}$. Note that $\Omega_{GW}^{\rm plt}(f)\propto\sqrt{\mu}\propto v_{\Phi}$, implying, the larger the symmetry breaking scale, the stronger the amplitude of GW. In our scenario, which features instead an intermediate matter-dominated epoch, the plateau breaks at a high frequency $f_{\rm dec}$, beyond which the spectrum falls as $\Omega_{GW}(f>f_{\rm dec})\sim f^{-1}$ (see Fig. 1). The analytical expression for this spectral break frequency $f_{\rm dec}$ can be computed as [57]
$f_{\rm dec}\simeq\sqrt{\frac{8z_{\rm eq}}{\alpha\Gamma G\mu}}\left(\frac{t_{\rm eq}}{t_{\rm dec}}\right)^{1/2}t_{0}^{-1}\,,$ (10)
where $z_{\rm eq}\simeq 3400$ is the red-shift at the standard matter-radiation equality taking place at time $t_{\rm eq}$, while $t_{\rm dec}$ is the time at which the scalar field $\Phi$ decays (matter domination ends). For the three benchmark cases in Fig. 1, the break of the GW spectrum occurs in the LISA [36], DECIGO [37] and ET [38] frequency bands, respectively.

### v2

@@ Abstract
Cosmic strings radiate detectable gravitational waves in models featuring high-scale symmetry breaking, e.g., high-scale leptogenesis. In this Letter, for the first time, we show that different flavored regimes of high-scale leptogenesis can be tested with the spectral features in cosmic string-radiated gravitational waves. This is possible if the scalar field that makes right-handed neutrinos massive is feebly coupled to the Standard Model Higgs. Each flavored regime, sensitive to low-energy neutrino experiments, leaves a marked imprint on the gravitational waves spectrum. A three-flavor and a two-flavor regime could be probed by a characteristic fall-off of the gravitational wave spectrum at the LISA-DECIGO-ET frequency bands with preceding scale-invariant amplitudes bounded from above and below. We present Gravitational Waves windows for Flavored Regimes of Leptogenesis (GWFRL) testable in the upcoming experiments. We also provide the first construction of a leptogenesis framework where a testable distinction of flavor regimes is possible without constraining the flavor structure of the theory.
@@
Lately, there have been efforts to obtain a direct possible signature of high-scale leptogenesis with primordial stochastic Gravitational Waves background (GWs) by focusing on the RHN sector. This is promising because, unlike electromagnetic radiation, GWs traverse the early universe unimpeded with unerring information about their origin. Therefore, no matter how high the leptogenesis scale is, it can be probed with GWs. The idea arises from contemplating the origin of RHN mass at a more fundamental level [14, 15, 16, 17]. Indeed, in Ref. [14] it has been argued that the RHNs become massive because of the spontaneous breaking of a $U(1)_{B-L}$ gauge invariance, which many Grand Unified Theories (GUT) might embed [18, 19, 20, 21]. Besides generating RHN mass, a gauged $U(1)_{B-L}$ breaking inevitably produces cosmic strings (here we consider only stable cosmic strings [22]) that radiate gravitational waves, which may be a signature of leptogenesis. Let us consider the $U(1)_{B-L}$ invariant RHN mass term $f_{N}\overline{N_{R}}\Phi N^{C}_{R}$, where $f_{N}$ is the Yukawa coupling, $N_{R}$ is the RHN field and $\Phi$ being the $B-L$ scalar with vacuum expectation value $v_{\Phi}$. After the $U(1)_{B-L}$ breaking, the RHNs obtain mass given by $M_{N}=f_{N}v_{\Phi}$, cosmic strings are formed [23, 24, 25] and radiate GWs with a scale-invariant spectrum at higher frequencies with an overall amplitude $\Omega_{\rm GW}\propto v_{\Phi}$ [26, 27]. While the scenario is interesting, note however, that the GW spectrum being dependent only on $v_{\Phi}$ is not so sensitive to $M_{N}$ because of the additional free parameter $f_{N}$. As such, a GW spectrum for $v_{\Phi}=10^{14}$ GeV may correspond to 3FL for $f_{N}=10^{-6}$ as well as 2FL for $f_{N}=10^{-4}$. Therefore, a more realistic GW signature of leptogenesis requires dependence on both parameters $v_{\Phi}$ and $f_{N}$, and hence also on $M_{N}$.
@@
In this Letter, we show that the $M_{N}$-dependent GW spectrum and a concrete link between GWs and high-scale leptogenesis can be achieved in the $U(1)_{B-L}$ model with a peculiar scalar field dynamics [28]. After the phase transition, $\Phi$ rolls down to its true vacuum and oscillates coherently to provide an early matter-domination when it is long-lived. The lifetime of the scalar field is determined by $v_{\Phi}$ and $f_{N}$, and therefore, $M_{N}$ controls the beginning and the end of the matter-domination. Differently from Ref. [14], the amplitude and the spectral features of the GWs radiated from cosmic strings are now imprinted by the $M_{N}$-dependent matter domination, thus carrying information on $M_{N}$, and therefore, on different regimes of flavored leptogenesis. In particular, the scale invariance in the GW spectrum at higher frequencies breaks at a frequency $f_{\rm dec}$ which is determined by the end of the matter domination at the temperature $T_{\rm dec}(M_{N})$ (see Fig. 1). We point out that such a scenario allows us to probe the RHN mass scale markedly, in agreement with all the relevant constraints on the parameter space (see Fig. 2). In particular, the prominent ones are: i) the bound from the recent observations of nHz GWs by Pulsar Timing Arrays (PTA) [29, 30, 31, 32, 33] which disfavor GWs from stable cosmic strings [34]; ii) the condition that the cosmic strings radiate predominantly GWs than particles [35, 36]; iii) the requirement that the $\Phi$ particles decay before the Big-Bang-Nucleosynthesis (BBN) [37, 38]; and iv) the scalar field $\Phi$ is long-lived. We highlight that if the break of the GW spectrum occurs in the LISA (DECIGO) frequency band, it is a signature (most likely signature) of a 3FL regime (see Fig. 3). On the other hand, a break of scale invariance in ET would correspond to both regimes, but with less signal strength for a 3FL. Remarkably, the amplitudes and the spectral break frequencies of the GW spectra predicted by our model are very well delimited and they are all within reach of planned gravitational wave experiments such as LISA, DECIGO, and ET [39, 40, 41]. Our results not only ensure testable GW signatures of different flavor regimes but also imply that the latter can be probed without constraining the flavor space of the leptogenesis Lagrangian.
@@
The $M_{N}$-dependent matter domination and constraints: The seesaw Lagrangian for neutrino masses and leptogenesis is given by
@@
where $f_{D}$ is the neutrino Dirac Yukawa coupling and $H(L)$ is the SM Higgs (lepton doublet). The fields $N_{R}$ and $L_{\alpha}$ have $B-L$ charge $-1$, whereas the same charges for $\Phi$ and $H$ are $2$ and 00, respectively. The scalar field dynamics is dictated by the finite temperature potential restoring the symmetry at higher temperatures given by [42, 43]
@@
where $D,E$ and $T_{0}$ are functions of $B-L$ gauge coupling $g^{\prime}$, the self-interaction coupling $\lambda$ and the vacuum expectation value $v_{\Phi}=m/\sqrt{\lambda}$, which is determined by the zero temperature potential $V(\Phi,0)=-\frac{m^{2}}{2}\Phi^{2}+\frac{\lambda}{4}\Phi^{4}$. The last term in Eq. (2) generates a potential barrier causing a secondary minimum at $\Phi\neq 0$, which becomes degenerate with the $\Phi=0$ one at $T=T_{c}$. The potential barrier vanishes at $T_{0}~{}(\lesssim T_{c})$, and the minimum at $\Phi=0$ becomes a maximum. The transition from $\Phi=0$ to $\Phi=v_{\Phi}$ can be treated as a second-order transition with an extremely quick disappearance of the potential barrier if the order parameter $\Phi_{c}/T_{c}\ll 1$. Such a smooth transition can be obtained approximately within the ballpark of $\lambda\simeq{g^{\prime}}^{3}$ and for ${g^{\prime}}\lesssim 10^{-2}$ [15, 44]. Once $\Phi$ rolls down, it oscillates around $v_{\Phi}$ behaving as matter [45, 46, 28]. Three important observations are in order. First, the $U(1)_{B-L}$ breaking scale can be as high as the GUT scale $\Lambda_{\rm GUT}$, implying $v_{\Phi}\lesssim 5\times 10^{15}~{}{\rm GeV}$. Second, the Hubble friction at $T_{c}$ should also be negligible for the rolling of the field down to $v_{\Phi}$; this can be achieved by considering $m_{\Phi}$ larger than the Hubble parameter $\mathcal{H}(T_{c})$. Third, RHNs become massive after the phase transition at $T=T_{c}$. Therefore, the scale of leptogenesis is bounded from above as $T_{\rm lepto}\sim M_{N}\lesssim T_{c}\simeq 2\sqrt{g^{\prime}}v_{\Phi}$.
@@
We aim to make $\Phi$ long-lived to induce a matter-domination epoch with a lifetime determined by $v_{\Phi}$ and $f_{N}$. In high-scale leptogenesis scenario, the coupling $f_{N}$ should be generically large to have $M_{N}$ comparable with $v_{\Phi}$. In such a case, if $m_{\Phi}>2M_{N}$, tree-level decay of $\Phi$ to RHN pairs ($\Phi\rightarrow NN$) would be too quick to make $\Phi$ long-lived. This restricts ourselves to the case $M_{N}>m_{\Phi}=\sqrt{2\lambda}v_{\Phi}$. Therefore, along with $M_{N}\lesssim T_{c}$, the $\Phi$ scalar is long-lived within the following range
$m_{\Phi}<M_{N}\lesssim T_{c}~{}~{}\implies~{}~{}\sqrt{2g^{\prime}}g^{\prime}<f_{N}\lesssim 2\sqrt{g^{\prime}}\,,$ (3)
@@
where we have assumed $\lambda={g^{\prime}}^{3}$. In this case, the mass of $\Phi$ is suppressed by a factor $\sqrt{g^{\prime}}$ compared to the $M_{Z^{\prime}}=\sqrt{2}g^{\prime}v_{\Phi}$, thus kinematically forbidding the tree-level decays of $\Phi$ to $Z^{\prime}$ pairs. Note interestingly that, even if defined to vanish at a very high scale, at lower energies, the seesaw Lagrangian in Eq. (1) can provide a nonvanishing effective coupling $\lambda_{\Phi H,{\rm 1-loop}}\sim f_{D}^{2}f_{N}^{2}/(2\pi^{2})$ at one-loop to trigger $\Phi\rightarrow HH$ decay  [47, 48, 28]. Therefore, despite $f_{N}$ being large, we can have a lower decay rate proportional to $\lambda_{\Phi H,\text{1-loop}}^{2}$. This decay involves both the RHN mass and the active neutrino mass $m_{\nu}$ via the Dirac Yukawa coupling $f_{D}=\sqrt{m_{\nu}f_{N}v_{\Phi}/v_{H}^{2}}$. In our discussion, we consider this radiative $\Phi\rightarrow HH$ decay with the assumption that $\lambda_{\Phi H,\text{1-loop}}\gg\lambda_{\Phi H,\text{tree}}$ (This the crucial theoretical identification in the $U(1)_{B-L}$ model, that makes GW spectrum sensitive to $M_{N}$). The decay rate is given by [47, 48, 28]
$\Gamma_{\Phi}^{HH}\simeq\Gamma_{0}\,\frac{f_{N}^{6}}{\lambda}\left(\frac{v_{\Phi}}{\rm 10^{13}~{}GeV}\right)^{2}\left(\frac{m_{\Phi}}{\rm 10^{8}~{}GeV}\right)\ln^{2}\left(\frac{\Lambda}{\tilde{\Lambda}}\right)\,,$ (4)
where the running of the one-loop effective coupling is determined by the UV cut-off scale, e.g., $\Lambda\simeq\Lambda_{\rm GUT}$ and the relevant physical scale $\tilde{\Lambda}$. We have $\Gamma_{0}=1.3\times 10^{-2}~{}{\rm GeV}$ assuming $m_{\nu}\simeq 0.01$ eV and the SM Higgs vacuum expectation value $v_{H}=174$ GeV.
@@
The evolution of the energy density of $\Phi$ and the duration of the matter-dominated epoch, which ends when $\Phi$ decays [28], can be tracked by solving the Friedmann equations (see Supplemental Material I for further details). We can compute the decay temperature of the scalar field by recursively solving the implicit equation $\Gamma_{\Phi}^{HH}(T_{\rm dec})\simeq\mathcal{H}(T_{\rm dec})$ taking $\tilde{\Lambda}\simeq T_{\rm dec}$ in Eq. (4). The scalar must decay before BBN, i.e., $T_{\rm dec}\gtrsim T_{\rm BBN}\sim 10~{}{\rm MeV}$. Moreover, because the scalar field is long-lived, it produces entropy which dilutes any pre-existing relic. Therefore, in the computation of baryon asymmetry, the entropy dilution must be considered, because in this framework, the produced entropy can be as large as $10^{6}$ (see Supplemental Material I). Finally, we mention that we can safely neglect the additional one-loop decay $\Phi\rightarrow f\bar{f}V$ (the two-body decay $\Phi\rightarrow f\bar{f}$ is suppressed due to chirality flip [49]), where $f$ and $V$ are SM fermions and vector bosons. Indeed, this decay dominates over the $\Phi\rightarrow HH$ only for $g^{\prime}\gtrsim 10^{-2}$ and smaller values of $v_{\Phi}$, which are outside our parameter space.
@@
Imprints of flavored leptogenesis on GWs from cosmic strings: Gravitational waves are radiated from cosmic string loops chopped off from the long strings resulting from the spontaneous breaking of the gauged $U(1)_{B-L}$ [26, 27]. Long strings are described by a correlation length $L=\sqrt{\mu/\rho_{\infty}}$, where $\rho_{\infty}$ is the long string energy density and $\mu$ is the string tension defined as [50]
@@
$\Omega_{\rm GW}=\sum_{k=1}^{k_{\rm max}}\frac{2k\mathcal{F}_{\alpha}G\mu^{2}\Gamma_{k}}{f\rho_{c}}\int_{t_{i}}^{t_{0}}\left[\frac{a(t)}{a(t_{0})}\right]^{5}n_{\omega}\left(t,l_{k}\right){\rm d}t\,,$ (6)
@@
where $\beta=2/3(1+\omega)$ with $\omega$ being the equation of state parameter of the universe, and $A_{\beta}=5.4$ ($A_{\beta}=0.39$) for radiation-dominated (matter-dominated) universe [56]. The quantity $\Gamma_{k}={\Gamma k^{-\delta}}/{\zeta(\delta)}$ quantifies the emitted power in $k$-th mode, with $\delta=4/3$ ($\delta=5/3$) for loops containing cusps (kinks) [57].
@@
Figure 1: The GW spectrum from the cosmic strings predicted by our model for four benchmark cases with $g^{\prime}=10^{-4}$: BP1 (2FL), $M_{N}=1.6\times 10^{9}~{}{\rm GeV}$ and $f_{N}=1.4\times 10^{-5}$; BP2 (3FL), $4.2\times 10^{8}~{}{\rm GeV}$ and $f_{N}=1.8\times 10^{-5}$; BP3 (3FL), $5.2\times 10^{6}~{}{\rm GeV}$ and $f_{N}=6.8\times 10^{-6}$; BP4 (3FL), $2.7\times 10^{8}~{}{\rm GeV}$ and $f_{N}=2.2\times 10^{-6}$. The points show the calculations of the GW plateau amplitude $\Omega_{\rm GW}^{\rm plt}$ and the break frequency $f_{\rm dec}$ according to Eq.s (8) and (9), respectively. BP1 and BP2 correspond to equal $\Omega_{\rm GW}^{\rm plt}$ but different $f_{\rm dec}$, while BP1 and BP4 to equal $f_{\rm dec}$ but different $\Omega_{\rm GW}^{\rm plt}$. BP3 represents the weakest allowed signal for $g^{\prime}=10^{-4}$ (see Fig. 2).
The integral in Eq. (6) is subjected to two Heaviside functions $\Theta(t_{i}-t_{\rm fric})\Theta(t_{i}-l_{c}/\alpha)$, each setting a cut-off at a very high frequency $f_{*}$, beyond which the GW spectrum falls as $f^{-1}$. A similar cut-off owing to inflation-diluted strings [58] is not so relevant to our context because we consider thermal leptogenesis. The quantity $t_{\rm fric}$ represents the time until which the motion of the string network is damped by friction [59], and $l_{c}$ is a critical length above which GW emission is dominant over particle production as shown by high-resolution numerical simulations [35, 36] (see also Ref. [60]). The critical length is $l_{c}\simeq\delta_{w}(\Gamma G\mu)^{-\gamma}$, where $\delta_{w}=(\sqrt{\lambda}v_{\Phi})^{-1}$ is the width of the string, and $\gamma=2$ ($\gamma=1$) for loops containing cusps (kinks). We shall neglect the cut-offs due to the kinks and friction as they are weaker than those from cusps. Here we deal with small values of $\lambda$ leading to “thick” strings, which are not conventional while studying GWs signatures from cosmic strings in BSM models [14]. In this case, the particle production cut-off in the GW spectrum is expected to be much stronger than the one usually considered for “thin” strings ($\lambda\simeq 1$). Thus, our scenario is also the first concrete BSM model that discusses the results of Ref.s [35, 36] for small $\lambda$. Note however that, we generalized the simulation results obtained for large couplings ($\lambda,g^{\prime}$) by replacing $\delta_{w}=(\sqrt{\lambda}v_{\Phi})^{-1}$ in $l_{c}$. But in principle, Eq.(11) of Ref. [35] would exhibit coupling dependence which is not trivial to deduce without numerical simulation. Therefore, the particle production cut-off discussed above should be considered conservative. Note however that the strings predicted in this model effectively evolve as thin strings because the condition $\delta_{w}\ll\mathcal{H}^{-1}$ is always satisfied (see Supplemental Material II). In the following, we present the results for the dominant $k=1$ mode which captures the qualitative features of our analysis.
Figure 2: The allowed parameter space (GWFRL window) in the $f_{\rm dec}$-$\Omega_{\rm GW}^{\rm plt}\,h^{2}$ plane for $g^{\prime}=10^{-4}$ with the leptogenesis scale $M_{N}$ color-coded. The different black lines display the constraints according to the discussed model requirements and the PTA measurements. The different points correspond to the benchmark cases displayed in Fig. 1. The orange solid line marks the transition between three-flavor and two-flavor leptogenesis regimes.
@@
In Fig. 1, we show the GW spectrum from Eq. (6) (for $k=1$) predicted by our model for four benchmark scenarios mentioned in the caption. Without any intermediate matter-dominated epoch, the GW spectrum has two main features: i) a low-frequency peak owing to the GW radiation from the loops that originate in the radiation epoch and decay in the standard matter epoch, and ii) an almost scale-invariant plateau at high frequencies as [51, 52, 61]
@@
$\Omega_{\rm GW}^{\rm plt}=\frac{128\pi\mathcal{F}_{\alpha}G\mu}{9\zeta(\delta)}\frac{A_{R}}{\epsilon_{R}}\Omega_{R}\left[(1+\epsilon_{R})^{3/2}-1\right]\,,$ (8)
@@
that arises from loop dynamics in the radiation-dominated epoch only. In Eq. (8), $\epsilon_{R}=\alpha/\Gamma G\mu\gg 1$, $A_{R}\equiv A_{\beta}\simeq 5.4$ and $\Omega_{R}\sim 9\times 10^{-5}$. Note that $\Omega_{\rm GW}^{\rm plt}\propto\sqrt{\mu}\propto v_{\Phi}$, implying the larger the symmetry breaking scale, the stronger the amplitude of GW. In our scenario, which features instead an intermediate matter-dominated epoch, the plateau breaks at a high frequency $f_{\rm dec}$, beyond which the spectrum falls as $\Omega_{\rm GW}(f>f_{\rm dec})\propto f^{-1}$ (see Fig. 1). The analytical expression for this spectral break frequency $f_{\rm dec}$ can be estimated as [62]
$f_{\rm dec}\simeq 0.45\,{\rm Hz}\left(\frac{10^{-12}}{G\mu}\right)^{1/2}\left(\frac{T_{\rm dec}}{\rm GeV}\right)\,.$ (9)